OpenStudy (anonymous):

limit x-> 0 of x/ [ abs(2x-1) - abs(2x+1) ] what is the limit?

OpenStudy (aravindg):

help mee

OpenStudy (nikvist):

what?

OpenStudy (aravindg):

i will post the question .can u say how to do?each step?

OpenStudy (nikvist):

\[\lim_{x\rightarrow 0}\frac{x}{|2x-1|-|2x+1|}=\lim_{x\rightarrow 0}\frac{x}{1-2x-(2x+1)}=\]\[=\lim_{x\rightarrow 0}\frac{x}{-4x}=-\frac{1}{4}\]